OpenStudy (anonymous):

HELP PLEASE ASAP. I really dont get it :((( 3 / 0.8 = 12 / b 3/4 = 21 / 2+d 4a / 9 = 100 / 40 / = Fraction bar.

OpenStudy (anonymous):

b=12 x 0.8/3 =3.2

OpenStudy (anonymous):

d=26

OpenStudy (anonymous):

a=45/8

OpenStudy (anonymous):

and b=3.2

OpenStudy (anonymous):

4a / 9 = 100 / 45 5 25 -5 -25

OpenStudy (anonymous):

I didnt mean to put 40

OpenStudy (anonymous):

5